# Rethinking the Atmospheric Scattering-driven Attention via Channel and Gamma Correction Priors for Low-Light Image Enhancement

### v1

@@ Abstract
Low-light image enhancement remains a critical challenge in computer vision, as does the lightweight design for edge devices with the computational burden for deep learning models. In this article, we introduce an extended version of Channel-Prior and Gamma-Estimation Network (CPGA-Net), termed CPGA-Net+, which incorporates an attention mechanism driven by a reformulated Atmospheric Scattering Model and effectively addresses both global and local image processing through Plug-in Attention with gamma correction. These innovations enable CPGA-Net+ to achieve superior performance on image enhancement tasks, surpassing lightweight state-of-the-art methods with high efficiency. Our results demonstrate the model’s effectiveness and show the potential applications in resource-constrained environments.
@@
Keywords Atmospheric Scattering Model, Attention Mechanism, Low-Light Image Enhancement, Gamma Correction, Channel Prior.
1 Introduction
@@
Low-light image capture, whether indoors or outdoors, poses significant challenges for accurate visual analysis. The limited light reflection often results in degraded image quality, including color inaccuracies and increased noise levels. These issues can significantly affect the performance and reliability of time-sensitive applications, such as transportation surveillance and Advanced Driver Assistance Systems (ADAS). Therefore, it is crucial to address these challenges to ensure the effective operation of systems under low-light conditions.
@@
The problem of low-light image enhancement is commonly addressed using two main methods: Histogram Equalization [1] and Retinex [2]. Histogram Equalization (HE) works by enhancing contrast through the redistribution of grayscale values. On the other hand, the Retinex theory divides the image into reflectance and illumination components to improve reflectance and overall image quality. Techniques such as Single Scale Retinex (SSR) [3] and Multi-Scale Retinex (MSR) [4] are particularly effective in preserving details and managing complex lighting conditions.
@@
With technological advancements, various deep learning-based methods have been proposed in recent years to enhance the quality of low-light images [5, 6, 7]. However, these methods often require substantial computational resources, which limits their practical application on real-world devices. Therefore, designing lightweight and efficient image enhancement techniques is crucial. In our previous study, we introduced the CPGA-Net [8], which combines Retinex theory with the Atmospheric Scattering Model (ATSM) [9] and utilizes gamma correction for both global and local processing; it highlights the importance of gamma correction in low-light image enhancement. CPGA-DIA [10] explores exposure correction and low-light enhancement issues through dynamic gamma adjustment, proving that gamma-correction prior is efficient for the enhancement process.
Building on these ideas with a theoretical-based structure and gamma-correction prior, we expanded CPGA-Net [8] by rethinking the theoretical equation as an attention mechanism and transitioning the gamma estimation module from global processing to local processing. Our main contributions are as follows:
  • Extended CPGA-Net: We propose an enhanced version of the Channel-Prior and Gamma-Estimation Neural Network (CPGA-Net) called CPGA-Net+. This model achieves state-of-the-art image quality and efficiency performance, making it a lightweight and practical solution for real-world applications, as shown in Fig. 1
  • Atmospheric Scattering-driven Attention: We introduce the Atmospheric Scattering Model as attention into the local network, named “Channel-Prior Block.” It improves the capture of channel prior information for better preservation and refinement of image structure and details.
  • Plug-in Attention with Gamma Correction: We integrate gamma correction into the local branch, transforming it into CPGA-Net-like attention modules. This maximizes the efficiency of environmental factors and enhances image quality through plug-in attention mechanisms.
@@
Figure 1: Performance Comparison with state-of-the-art approaches. (a) Comparison of PSNR vs. Number of Parameters on LOLv1; (b) Comparison of mean NIQE vs. Number of Parameters on unpaired data.
2 Related Works
@@
Our work enhances the original purely convolutional architecture by integrating an attention mechanism while preserving the benefits of a lightweight and efficient design. This advancement is particularly well-suited for low-light image enhancement tasks. In this section, we will conduct a comprehensive literature review on leveraging a deep learning-based approach in low-light image enhancement and explore developments in lightweight model architectures.
2.1 Deep Learning-Based Low-Light Image Enhancement
@@
With the continuous development of low-light image enhancement (LLIE), Retinex theory has increasingly demonstrated its potential in conjunction with deep learning techniques. Several methods based on this approach address low-light environments. For instance, while LIME [11] differs from directly decomposing images according to Retinex theory, it primarily relies on estimating the illumination map of low-light images for enhancement. RetinexNet [5] and KinD [7] decompose images into reflectance and illumination components during decomposition. In the adjustment phase, they adjust the illumination component’s brightness and denoise the reflectance component, ultimately merging them based on the theory to restore natural images. EnlightenGAN [6] proposes an unsupervised Generative Adversarial Network (GAN) that trains without paired low/normal light images. It introduces a global-local discriminator structure, self-regularized perceptual loss fusion, and attention mechanisms to enhance image quality. LLFlow [12] presents a flow-based approach for low-light image enhancement by modeling distributions of normally exposed images. It improves traditional methods by using an illumination-invariant color map as the prior distribution rather than a Gaussian distribution. The process features an encoder to extract stable color attributes and an invertible network to map low-light images to normally exposed image distributions, aiming for improved enhancement performance.
2.2 Lightweight Low-Light Image Enhancement
@@
In the context of low-light image enhancement, developing lightweight methods is crucial for practical deployment, often requiring sophisticated techniques to achieve both efficiency and effectiveness. For example, Zero-DCE [13] replaces the direct image enhancement process with a curve-fitting approach, introducing a series of non-reference loss functions that reduce the computational burden, achieving an efficient and lightweight design. RUAS [14] builds upon Retinex theory by proposing a Retinex-inspired model that leverages prior information from low-light images, combined with a distillation unit-based search architecture and a cooperative bilevel search strategy, maintaining high performance while achieving a lightweight design. IAT [15] decomposes the task into local and global processing components. The local branch leverages a convolution-based Transformer to perform image restoration and enhancement. In contrast, the global branch utilizes global priors, including color transformation matrices and gamma correction, to apply global adjustments across different exposure conditions, thereby attaining efficient and lightweight performance improvements. PairLIE [16] deviates from the traditional Retinex approach of directly decomposing images; instead, it removes noise through a self-supervised mechanism before decomposition. Training on low-light images of the same scene with different exposures better learns features. Finally, it merges them using a simple convolutional network to achieve a lightweight design. Inspired by Retinex theory and ISP frameworks, FLIGHT-Net [17] features Scene Dependent Illumination Adjustment (SDIA) for illumination and gain processing and Global ISP Network Block (GISP) for compact color correction and denoising. This design optimizes for both efficiency and lightweight operation.
3 Methodology
@@
In this section, we delve into the reconstruction of ideal images by global and local concepts in image processing, leveraging advanced deep learning techniques. The discussion will commence with the theoretical underpinnings and motivations for developing CPGA-Net+, followed by a rigorous exposition of the network’s architecture and implementation, which can be separated as Atmospheric Scattering-driven Attention and Plug-in Attention with Gamma Correction, as shown in Fig. 2.
@@
3.1 The Connection Between Theoretical Equations and Low-light Image Enhancement
Guo et al. [11] figured out that the Retinex theory [2] and the Atmospheric Scattering Model (ATSM) [9] have a strong relationship that can represent each other by reformulating both equations. Retinex theory assumes that the received image can be decomposed into illumination and reflectance components, as shown in Eq. (1):
$I(i,j)=L(i,j)\cdot R(i,j)$ (1)
for each pixel location at $(i,j)$, where $R(i,j)$ represents the reflectance component, $I(i,j)$ is the perceived image, and $L(i,j)$ denotes the illumination component.
On the other hand, the Atmospheric Scattering Model is defined as:
$J(x)=\frac{I(x)-A(x)}{t(x)}+A(x)$ (2)
where $I(x)$ represents the input image, $J(x)$ represents the haze-free image, $t(x)$ represents the atmospheric transmission, and $A(x)$ represents the intensity of atmospheric light.
Based on Dong et al. [18], the low light image $I(x)$ can be seen as $1-L(x)$, and $J(x)$ can be seen as $1-R(x)$. Thus, we can rewrite it into the following form:
$R(x)=\frac{L(x)-\tilde{A}(x)}{t(x)}+\tilde{A}(x)$ (3)
where $\tilde{A}(x)=1-A(x)$. Notably, if $\tilde{A}=0$, the equation aligns with the Retinex theory.
In our previous work, CPGA-Net [8], we successfully utilized these theoretical equations in deep learning form. In this article, we extended the idea and proposed an attention mechanism called the Channel-Prior block (CP block), which modularizes the relationship of both equations into a systematic form. We restructured the module with convolutions and a ResBlock while fusing features with the original RGB channels at each step to streamline the module into an attention-block design. This helps in activating the feature map to align with the original channels. In CPGA-Net, three-channel priors are selected as the input for $t$ estimation: the Bright Channel Prior (BCP), the Dark Channel Prior (DCP), and the luminance channel (Y component from the YCbCr color space). They can be defined respectively as:
@@ 3.1 The Connection Between Theoretical Equations and Low-light Image Enhancement
where $I^{c}$ represents the color channel $c$ of the input image $I$. These common features represent the brightness variation in different environments and are widely used in traditional methods [19]. The combination of channel priors shows sensitivity to brightness, which is an important clue in representing the atmospheric transmission $t$ to guide the enhancement, as shown in Fig. 3.
Figure 3: The comparison between the brightness representation. From top to bottom are the original image, edge detection of the luminance channel $(I^{\text{luminance}})$, and edge detection of the channel-priors, which is $\max(I^{\text{bright}},I^{\text{dark}},I^{\text{luminance}})$.
Additionally, to transform it into a high-dimensional processing module rather than maintaining the original channels, we simplified the luminance channel into a more basic representation—specifically, the mean of the channels, akin to the concept of intensity (I component in the HSI color space), which also indicates image brightness. The channel priors are depicted as:
@@
where $F$ is the input feature map, and $F^{CP}$ is the channel prior features. If $c$ consists of RGB channels, $F$ will be equal to $I$. As a result, the channel priors are simplified to the input channels’ max, min, and mean, making the attention module more responsive to the overall brightness control of the image.
For the $\tilde{A}$ estimation, which captures detailed features and reconstructs the image, we redesigned it as a mini-U-Net-based architecture with encoder and decoder pathways connected by skip connections. This design effectively captures and preserves spatial information at multiple scales, enhancing the model’s ability to reconstruct fine details in the image—a technique commonly used in low-light image enhancement, such as [7, 6]. Due to considerations for lightweight efficiency, we only downsample the input once, reducing computational complexity while maintaining adequate feature extraction capabilities. This approach ensures that the model remains efficient and suitable for real-time applications or scenarios with limited computational resources without significantly compromising the quality of the reconstructed image.
After obtaining the estimates of $t$ and $\tilde{A}$, we can reconstruct our features using Eq. (3), which serves as an attention module sensitive to brightness variations in the scene. This leads to the proposed Atmospheric Scattering-driven Attention, formulated as follows:
$R_{\text{att}}(F)=\frac{L^{\prime}(F)-\tilde{A}(F)}{t(F,F^{CP})}+\tilde{A}(F)$ (8)
where $L^{{}^{\prime}}$ is the mapped input tensor with a matching channel for formula calculation, and the structure is shown in Fig. 4.
Figure 4: The block diagram of the Channel-Prior Block (CP block).
3.2 Relinking the Gamma Correction to Global-Local Processing
@@
where $\gamma$ is the gamma value controlling the degree of correction, enhancing the input image $r$ to produce the output image $s$. These approaches combined an independent branch of regression with the enhancement model to better estimate gamma values. Moreover, the complexity of gamma value estimation can complicate training goals and make the process prone to divergence. Taking these aspects into account, an Intersection-Aware Adaptive Fusion Module (IAAF) was proposed, as shown in Fig. 5 and Eq. (10):
$\hat{R}=\text{IAAF}(R^{\gamma},R)=(R\cup R^{\gamma})-(R\cap R^{\gamma})\approx R+R^{\gamma}-\cap(R,R^{\gamma})$ (10)
where the enhanced image $\hat{R}$ is created by combining $R$ and $R^{\gamma}$ while removing any overlapping elements and $\cap(R,R^{\gamma})$ represents the estimation of the intersection for finding the features with similarity across the gamma-corrected and original image.
Figure 5: Intersection-Aware Adaptive Fusion Module (IAAF) [8].
First, based on the findings from [20, 21], we know that downsampling the image resolution by a limited scale does not significantly impact performance while greatly reducing computational cost. Therefore, we redesigned the process of gamma value estimation, as shown in Fig. 2. We reduced the feature resolution in the first ResCBAM block by adjusting the stride from 1 to 2. This allows subsequent computations to operate on half the image size, incorporating a residual from the preprocessing layer of gamma correction with downsampling. It enhances the learnability of the fusion strategy between gamma correction and the local branch.
Furthermore, considering the standard gamma correction operation, as shown in Eq. (9). In our attention mechanism, we adapt this operation to function within the Channel-Prior Block by incorporating it directly into the attention process, allowing it to act as a global feature that estimates the optimal gamma correction for each feature channel, rather than applying gamma correction only at the final step as commonly done in [8, 15]. The modified attention output can then be expressed as:
@@
where $\hat{R}_{\text{att}}$ is the attention feature built on the gamma-corrected feature $R_{\text{att}}^{\gamma}$ and the original feature $R_{\text{att}}$, respectively. Moreover, we add another residual to this attention, distinguishing it from the reconstruction applied in Eq. (10), making it an auxiliary attention to support the residual features.
Building on insights from [10], we understand that the adaptive gamma value serves as an environmental factor representing the overall illumination conditions, which can vary across different scenes. Consequently, this attention mechanism ensures that the network focuses on regions within the broader image where gamma correction yields the most significant enhancement. This approach is called “Plug-in Attention with Gamma Correction” for the “CPGA block.”
4 Experiment Results
This section analyzes our approach using several state-of-the-art methods on benchmark datasets.
4.1 Implementation Details
Our experiments are conducted and evaluated with an NVIDIA GeForce RTX 3090 GPU. For the preprocessing of training, we cropped the image into $256\times 256$ pixels with random translation. Adam optimization with a learning rate of $10^{-3}$ (LOLv1 & LOLv2 Synthetic) and a learning rate of $10^{-4}$ (LOLv2 Real) was set for the training, with 600 epochs by Cosine Annealing scheduler and the last epoch set as 67 for a cycle.
4.2 Loss Functions
Four loss functions are used to supervise our approach, including L1 loss, perceptual loss, HDR loss, and SSIM loss. The L1 loss function, a commonly use loss function that performs better in image enhancement and restoration, is defined as:
$L_{1}=\|\hat{Y}-Y^{GT}\|_{1}$ (12)
where $\hat{Y}$ is the output and $Y^{GT}$ is the ground truth.
Perceptual loss [22] is a loss function commonly used in image restoration, style transfer, and image generation. It emphasizes capturing high-level features and structures that closely resemble human perception. The loss can be expressed as:
$L_{per}=\|\Psi(\hat{Y})-\Psi(Y^{GT})\|_{2}^{2}$ (13)

### v2

@@ Abstract
Enhancing low-light images remains a critical challenge in computer vision, as does designing lightweight models for edge devices that can handle the computational demands of deep learning. In this article, we introduce an extended version of the Channel-Prior and Gamma-Estimation Network (CPGA-Net), termed CPGA-Net+, which incorporates an attention mechanism driven by a reformulated Atmospheric Scattering Model and effectively addresses both global and local image processing through Plug-in Attention with gamma correction. These innovations enable CPGA-Net+ to achieve superior performance on image enhancement tasks for supervised and unsupervised learning, surpassing lightweight state-of-the-art methods with high efficiency. Furthermore, we provide a theoretical analysis showing that our approach inherently decomposes the enhancement process into restoration and lightening stages, aligning with the fundamental image degradation model. To further optimize efficiency, we introduce a block simplification technique that reduces computational costs by more than two-thirds. Experimental results validate the effectiveness of CPGA-Net+ and highlight its potential for applications in resource-constrained environments.
Index Terms:
@@
Atmospheric Scattering Model, Low-Light Image Enhancement, Gamma Correction, Channel Prior, Explainable AI.
I Introduction
@@
Low-light image capture, whether indoors or outdoors, poses significant challenges for accurate visual analysis. The limited light reflection often results in degraded image quality, including color inaccuracies and increased noise levels. These issues can significantly affect the performance and reliability of light-sensitive applications, such as transportation surveillance and Advanced Driver Assistance Systems. Therefore, it is crucial to address these challenges to ensure the effective operation of systems under low-light conditions.
@@
The problems of low-light image enhancement (LLIE) are commonly addressed using two main methods: Histogram Equalization [1] and Retinex [2]. Histogram Equalization works by enhancing contrast through the redistribution of grayscale values. On the other hand, the Retinex theory divides the image into reflectance and illumination components to improve reflectance and overall image quality. Techniques such as Single Scale Retinex [3] and Multi-Scale Retinex [4] are particularly effective in preserving details and managing complex lighting conditions.
@@
With technological advancements, various deep learning-based methods [5, 6, 7] have been proposed to enhance the quality of low-light images. However, these methods often require substantial computational resources, which limits their practical application on real-world devices. Therefore, designing lightweight and efficient image enhancement techniques is crucial. In our previous study, we introduced the CPGA-Net [8], which combines Retinex theory with the Atmospheric Scattering Model (ATSM) [9] and utilizes gamma correction for both global and local processing; it highlights the importance of gamma correction in LLIE. CPGA-DIA [10] explores exposure correction and low-light enhancement issues through dynamic gamma adjustment, showing that gamma correction can be efficient and effective for the enhancement process even in deep learning frameworks.
@@
Figure 1: Performance Comparison with SOTA approaches. (a) Comparison of PSNR vs. Number of Parameters on LOLv1; (b) Comparison of mean NIQE vs. Number of Parameters on unpaired data.
Building on these ideas with a theoretical-based structure and gamma-correction prior, we expanded CPGA-Net by rethinking the theoretical equation as an attention mechanism and transitioning the gamma estimation module from global processing to local processing. Our main contributions are as follows:
  • Extended CPGA-Net: We propose an enhanced version of the Channel-Prior and Gamma-Estimation Neural Network (CPGA-Net) called CPGA-Net+. This model achieves state-of-the-art (SOTA) image quality and efficiency performance for both supervised and unsupervised learning, making it a lightweight and practical solution for real-world applications, as shown in Fig. 1.
  • Theoretical-based Attention for Illumination: We modularized the Atmospheric Scattering Model into a block design and incorporated gamma correction into the local branch. This architecture modification significantly improves image structure and detail, maximizing the efficiency of prior knowledge for illumination to strengthen the overall image quality.
  • See Through the Mechanism Behind the Image Enhancement: We evaluate our approach through the degradation formula and optimization, demonstrating that the optimal solutions align with the output features from the neural network, ensuring strong interpretability and robustness.
  • Block Design Simplification for Lightweight Implementation: We can directly remove the local branch of our network during inference, which induces a slight performance drop but significantly improves efficiency, reducing the number of parameters and FLOPS by about 66.8% and 77.1%, respectively.
II Related Work
@@
Our work enhances the original purely convolutional architecture by integrating an attention mechanism while preserving the benefits of a lightweight and efficient design. This advancement is particularly well-suited for LLIE tasks. This section will conduct a comprehensive literature review on leveraging a deep learning-based approach in LLIE and explore developments in lightweight model architectures.
II-A Deep Learning-Based LLIE
@@
With the continuous development of LLIE, Retinex theory has increasingly demonstrated its potential in conjunction with deep learning techniques. Several methods based on this approach address low-light environments. For instance, while LIME [11] differs from directly decomposing images according to Retinex theory, it primarily relies on estimating the illumination map of low-light images for enhancement. RetinexNet [5] and KinD [7] decompose images into reflectance and illumination components during decomposition. In the adjustment phase, they adjust the illumination component’s brightness and denoise the reflectance component, ultimately merging them based on the theory to restore natural images. EnlightenGAN [6] proposes an unsupervised Generative Adversarial Network (GAN) that can be trained without paired low/normal light images. It introduces a global-local discriminator structure, self-regularized perceptual loss fusion, and attention mechanisms to enhance image quality. LLFlow [12] presents a flow-based approach for LLIE by modeling distributions of normally exposed images. It improves traditional methods by using an illumination-invariant color map as the prior distribution rather than a Gaussian distribution. The process features an encoder to extract stable color attributes and an invertible network to map low-light images to normally exposed image distributions, aiming for improved enhancement performance.
II-B Lightweight LLIE
@@
In the context of LLIE, developing lightweight methods is crucial for practical deployment, often requiring sophisticated techniques to achieve both efficiency and effectiveness. For example, Zero-DCE [13] replaces the direct image enhancement process with a curve-fitting approach, introducing a series of reference-free loss functions that reduce the computational burden, achieving an efficient and lightweight design. RUAS [14] builds upon Retinex theory by proposing a Retinex-inspired model that leverages prior information from low-light images, combined with a distillation unit-based search architecture and a cooperative bilevel search strategy, maintaining high performance while achieving a lightweight design. IAT [15] decomposes the task into local and global processing components. The local branch leverages a convolution-based Transformer to perform image restoration and enhancement. In contrast, the global branch utilizes global priors, including color transformation matrices and gamma correction, to apply global adjustments across different exposure conditions, thereby attaining efficient and lightweight performance improvements. PairLIE [16] deviates from the traditional Retinex approach of directly decomposing images; instead, it removes noise through a self-supervised mechanism before decomposition. It shows that training on low-light images of the same scene with different exposures better learns features. Finally, it merges them using a simple convolutional network to achieve a lightweight design. Inspired by Retinex theory and ISP (Image Signal Processor) frameworks, FLIGHT-Net [17] features Scene Dependent Illumination Adjustment for illumination and gain processing and Global ISP Network Block for compact color correction and denoising. This design optimizes for both efficiency and lightweight operation.
II-C Insights and Innovations
Our proposed method builds on two key insights from prior work: theory-driven attention mechanisms and lightweight design principles in LLIE. Our approach leverages the theoretical basis of perceiving visual information through air turbulence—a principle effectively utilized in Retinex-based [5, 6, 7, 11, 12, 14, 16] and ATSM methods [8, 10, 18]. We incorporate this concept into a streamlined attention module that selectively enhances key features, guided by prior knowledge [8, 13, 19], to improve detail preservation and contrast. Furthermore, following [8, 10, 15], gamma correction is integrated into our model and has extended its application from global to local processing to maximize adaptability via environmental characteristics. These attention modules and formulae turn our model apart by integrating theoretical principles with practical design, delivering high-quality enhancement while preserving a lightweight architecture.
III Methodology
@@
In this section, we delve into the reconstruction of ideal images by global and local concepts in image processing, leveraging advanced deep learning techniques. The discussion will commence with the theoretical underpinnings and motivations for developing CPGA-Net+, followed by an exposition of the network’s architecture and implementation, which can be separated as Atmospheric Scattering-driven Attention and Plug-in Attention with Gamma Correction, as shown in Fig. 2.
@@
III-A The Connection Between Theoretical Equations and Low-light Image Enhancement
Guo et al. [11] figured out that the Retinex theory [2] and the ATSM [9] have a strong relationship that can represent each other by reformulating their equations. Retinex theory assumes that the received image can be decomposed into illumination and reflectance components, as shown in Eq. (1):
$S=i\cdot R$ (1)
where $S$ is the perceived image, $i$ denotes the illumination component, and $R$ represents the reflectance component.
On the other hand, the Atmospheric Scattering Model for haze removal is defined as:
$I=tJ+(1-t)A$ (2)
where $I$ represents the input image, $J$ represents the haze-free image, $t$ represents the atmospheric transmission, and $A$ represents the intensity of atmospheric light.
Based on Dong et al. [18], the low light image $I$ can be seen as $1-L$, where $L$ represents the low-light image, and $J$ can be seen as $1-R$, where $R$ reflects the important characteristics of the input image. The above substitutions are performed under the normalized pixel values in [0, 1]. With these substitutions, we rewrite Eq. (2) into the following form:
$R=\tilde{t}L+(1-\tilde{t})\tilde{A}$ (3)
where $\tilde{A}=1-A$ and $\tilde{t}=1/t$. The model Eq. (3) described is the cornerstone of our neural network design. Part of the information of the reflectance $R$ comes from the known image $L$, part of it comes from an unknown image $\tilde{A}$, and the proportion sum of their contributions is limited to 1. When $L$ is very dark or noisy (the scene information is less reliable), the contribution of $L$ is lowered, and the contribution of $\tilde{A}$ is increased; when $L$ is relatively bright and less noisy (the scene information is more reliable), the contribution of $L$ is increased, and the contribution of $\tilde{A}$ is reduced. So, $\tilde{t}$ should reflect the intensity level of $L$ in some way.
This reformulation reveals an alternative imaging perspective, where $L$ is linked to the characteristic of atmospheric light, which predominantly includes environmental interference, $\tilde{A}$ corresponds to the unknown noise-free image, and $R$ corresponds to the reflectance in a linear relationship between $L$ and $\tilde{A}$. The underlying mechanism of this formulation matches the phenomenon of our atmospheric scattering-based approach and will be discussed in Section III-D.
In our previous work, CPGA-Net [8], we successfully utilized these theoretical equations in deep learning form. In this article, we extended the idea and proposed an attention mechanism called the Channel-Prior block (CP block), which modularizes the relationship of both equations into a systematic form. We restructured the module with convolutions and a ResBlock while fusing features with the original RGB channels at each step to streamline the module into an attention-block design. This helps in activating the feature map to align with the original channels. In CPGA-Net, three-channel priors are selected as the input for $\tilde{t}$ estimation: the Bright Channel Prior (BCP), the Dark Channel Prior (DCP), and the luminance channel (Y component from the YCbCr color space). They can be defined respectively as:
@@ III-A The Connection Between Theoretical Equations and Low-light Image Enhancement
where $I^{c}$ represents the color channel $c$ of the input image $I$. These common features represent the brightness variation in different environments and are widely used in traditional methods [19]. The combination of channel priors shows sensitivity to contrast, which is an important clue in representing the atmospheric transmittance $\tilde{t}$ to guide the enhancement, as shown in Fig. 3.
Figure 3: The comparison between the brightness representation. From top to bottom are the original image, edge detection of the luminance channel $I^{\text{luminance}}$, and edge detection of the channel-prior which is $\max(I^{\text{bright}},I^{\text{dark}},I^{\text{luminance}})$.
Additionally, to transform the priors into a high-dimensional processing module rather than maintaining the original channels, we simplified the luminance channel into a more basic representation—specifically, the mean of the channels, akin to the concept of intensity (such as the I component in the HSI color space), which also indicates image brightness. Therefore, the channel priors are depicted as:
@@
where $f$ is the input feature map, and $F^{\text{CP}}$ denotes the channel prior features. If $c$ consists of RGB channels, $f$ will equal the $I$ that appears in Eq. (4), (5), and (6) As a result, the channel priors are simplified to the input channel’s max, min, and mean, making the attention module more responsive to the overall contrast control of the image.
For the $\tilde{A}$ estimation, which captures detailed features and reconstructs the image, we redesigned it as a mini-U-Net-based architecture with encoder and decoder pathways connected by skip connections. This design effectively captures and preserves spatial information at multiple scales, enhancing the model’s ability to reconstruct fine details in the image—a technique commonly used in LLIE, such as [5, 6, 7]. Due to considerations for lightweight efficiency, we only downsample the input once, reducing computational complexity while maintaining adequate feature extraction capabilities. This approach ensures that the model remains efficient and suitable for real-time applications or scenarios with limited computational resources without significantly compromising the quality of the reconstructed image.
After obtaining the estimates of $\tilde{t}$ and $\tilde{A}$, we can reconstruct our features using Eq. (3), which serves as an attention module sensitive to brightness variations in the scene. This leads to the proposed Atmospheric Scattering-driven Attention, formulated as follows:
$R_{\text{att}}(f)=\tilde{t}(f,F^{\text{CP}})L^{\prime}(f)+\big{[}1-\tilde{t}(f,F^{\text{CP}})\big{]}\tilde{A}(f)$ (8)
where $L^{{}^{\prime}}$ is the mapped input tensor with a matching channel for formula calculation, $\tilde{t}(f,F^{\text{CP}})$ indicates the derived transmittance with input feature map $f$ and channel-prior features $F^{\text{CP}}$, and the structure is shown in Fig. 4.
Figure 4: The block diagram of the Channel-Prior Block (CP block), where c denotes a concatenation operation.
III-B Relinking the Gamma Correction to Global-Local Processing
@@
where $\gamma$ is the gamma value controlling the degree of correction, enhancing the input image $r$ to produce the output image $s$. These approaches combined an independent branch of regression with the enhancement model to better estimate gamma values. Moreover, the complexity of gamma value estimation can complicate training goals and make the process prone to divergence. Taking these aspects into account, we proposed an IAAF (Intersection-Aware Adaptive Fusion) module, as shown in Fig. 5 and Eq. (10):
$\begin{split}\hat{R}&=\text{IAAF}(R^{\gamma},R)=(R\cup R^{\gamma})-(R\cap R^{\gamma})\\ &\approx R+R^{\gamma}-\cap(R,R^{\gamma})\end{split}$ (10)
where the enhanced image $\hat{R}$ is created by combining $R$ and $R^{\gamma}$ while removing any overlapping elements and $\cap(R,R^{\gamma})$ represents the intersection estimation for finding similar features across gamma-corrected and uncorrected images.
Figure 5: A block diagram of the IAAF (Intersection-Aware Adaptive Fusion) module [8].
First, based on the findings from [20, 21], we know that downsampling the image resolution by a limited scale does not significantly impact performance while greatly reducing computational cost. Therefore, we redesigned the process of gamma value estimation, as shown in Fig. 2. We reduced the feature resolution in the first ResCBAM block by adjusting the stride from 1 to 2. This adjustment enables subsequent computations to process half-sized images while leveraging residual and downsampling connections. This design improves the learnability of the IAAF module, facilitating more effective integration of global and local processing.
Furthermore, we adapt this operation to function within the Channel-Prior Block by incorporating it directly into the attention process, allowing it to act as a global feature that estimates the optimal gamma-correction value for each feature channel rather than applying gamma correction only at the final step as commonly done in [8, 15]. The modified attention output can then be expressed as:
@@
where $\hat{R}_{\text{att}}$ is the attention feature built on the gamma corrected feature $R_{\text{att}}^{\gamma}$ and the uncorrected feature $R_{\text{att}}$. Moreover, we add another residual to this attention in Eq. (11), distinguishing it from the reconstruction applied in Eq. (10), making it an auxiliary attention to support the residual features.
Building on insights from [10], we understand that the adaptive gamma value serves as an environmental factor representing the overall illumination conditions, which can vary across different scenes. Consequently, this attention module ensures that the network focuses on regions within the broader image where gamma correction yields the most significant enhancement. This approach is called “Plug-in Attention with Gamma Correction,” and the CP block with IAAF is named “CPGA block.”
III-C Loss Functions
In supervised learning, we use four loss functions to guide our approach: L1 loss, perceptual loss, HDR L1 loss, and SSIM loss. We utilize HDR L1 loss, perceptual loss, and total variation loss for unsupervised learning, with the target set as the histogram-equalized low-light image.
The L1 loss function, a commonly use loss function that performs better in image enhancement and restoration, is defined as:
$L_{1}=\|\hat{Y}-Y^{\text{GT}}\|_{1}$ (12)
where $\hat{Y}$ is the output and $Y^{\text{GT}}$ is the ground truth.
Perceptual loss [22] is commonly used in image restoration, style transfer, and generation. It emphasizes capturing high-level features and structures that closely resemble human perception. The loss can be expressed as:
$L_{per}=\|\Psi(\hat{Y})-\Psi(Y^{\text{GT}})\|_{2}^{2}$ (13)